OpenStudy (deadshot):

Which of the following graphs represents a vertical shift down 1 of the parent graph of f(x)=sinx?

OpenStudy (kira_yamato):

f(x) = sinx - 1

OpenStudy (kira_yamato):

Remember, y = f(x) --> y - a = f(x) means a translation of a units in the positive y-axis direction

OpenStudy (deadshot):

Thanks!